In preparation for the upcoming school year, a teacher looks at raw test scores on the statewide standardized test for the students in her class. Instead of looking at the scores relative to the norms in the state, the teacher wants to understand the scores relative to the students who will be in the class. To do so, she decides to convert the test scores into z-scores relative to the mean and standard deviation of the students in the class.

The mean test score in her upcoming class is 49, and the standard deviation is 20.7.

The teacher now wants to identify those students who may need extra help. She decides to look at students who have z-scores below z = -2.00

Identify the test score corresponding to a z-score of below z = -2.00. Round to the nearest whole number
Answer #1


= 49

= 20.7

Z = - 2.00 = (X - 49)/20.7

So,

X = 49 - (2.00 X 20.7)

   = 49 - 41.4

   = 7.6 = 8 (Round to nearest whole number)

So,

Answer is:

8